# RapidPoseTriangulation: Multi-view Multi-person Whole-body Human Pose Triangulation in a Millisecond

### v1

@@ 1 Introduction
The ability to accurately detect and track human body positions and movements is a very important task in many advanced computer vision applications. From enhancing virtual reality experiences to improving the collaboration with humans in robotic systems, precise human pose estimation plays a crucial role in bridging the gap between digital systems and the physical world.
@@
However, marker-less pose estimation presents its own set of challenges. Occlusions, varying lighting conditions, and the complexity of human movements can all impact the accuracy and reliability of these systems. To address these issues, multi-view approaches have become more common. By capturing the scene from multiple angles simultaneously, these systems can overcome many of the limitations found in single-view methods.
Recent advances in deep learning have dramatically improved the accuracy of pose estimation from individual camera views. However, the task of efficiently combining these multiple 2D estimates into accurate 3D poses, particularly for multiple people in real-time scenarios, remains an active area of research. Furthermore, as applications become more advanced, there is an increasing demand for a more detailed whole-body pose estimation, including facial expressions and especially hand gestures (see Figure 1).
@@ 2 Related Work
Regarding the generalization of learning-based approaches to new datasets, a direct transfer is typically not evaluated. However, VoxelPose, Faster-VoxelPose, MvP, and TEMPO have implemented a finetuning strategy utilizing synthetic data, though only the first two have also published the source-code for this. This method involves randomly positioning 3D poses from an alternative dataset within 3D space and subsequently back-projecting them onto the camera perspectives of the target setup. These resulting 2D poses, after undergoing various augmentations, are then employed to learn the 3D reconstruction process. In the closed-source approach of CloseMoCap [16] this concept was further improved by using a larger 3D pose dataset and additional augmentations. Algorithmic approaches generally evaluated their transferability to a limited number of other datasets, a process that is considerably more straightforward as it doesn’t require any training.
@@ 3.1 Algorithm
Similar to most other approaches, the algorithm can be split into two stages as well, with the first one predicting the 2D poses for each image. For this any 2D pose estimator can be integrated, here RTMPose [12], similar as in VoxelKeypointFusion, was used. The second stage can be split into the following steps:
@@
In step (1) for each 2D pose, all possible pairs with poses from other views are created, since no association to real persons is known yet. In the next step (2), the pairs are filtered using the previous 3D poses, if available. For this, the 3D poses from the last frame are projected into the 2D views and a pixel-based distance threshold is used. The idea is to reduce the total number of pairs that need to be triangulated. In a pair of 2D persons in two different views, either both parts can match to one of the reprojected 3D poses, only one part is matched, or none of them. If both pairs match, it is likely a valid pair belonging to that 3D person, and it is kept. If only one part matches, a match is unlikely, because the 3D pose should be matchable in both views, so the second part is likely a pose from a different 3D person, and the pair is dropped. If none is matching, it is likely that a new 3D person is present in the current frame, and the pair is kept. In step (3) the core joints are extracted, which are the shoulders, hips, elbows, wrists, knees, and ankles. They are enough to associate the following 3D proposals to a person, and this reduces the computational complexity in the next steps. In step (4) each pair is triangulated to a 3D proposal (see Figure 2(b)). In the following step (5) clearly invalid proposals outside the observed room are dropped. In step (6) the remaining proposals are reprojected into the 2D views. Step (7) calculates a pixel-distance-based reprojection error to the pair’s original 2D poses, and also prunes clearly invalid limbs. In step (8) pairs/proposals with a large error are dropped, as they are considered invalid matches. In step (9) the remaining 3D proposals are grouped in 3D space. Each of the 2D view pairs created a 3D proposal, and if they are close in 3D space, they likely belong to the same person, and are therefore grouped together. Grouping in 3D makes use of physical constraints, which allows a simpler and faster implementation. In step (10) the remaining pairs, which normally are less than in step (4), are triangulated again, but this time with all joints. In step (11) for each group the calculated 3D proposals are merged into a single 3D pose. To filter outliers, first, for each joint an average location is calculated, and then for each proposal, the distance of this joint to the average is calculated. If the distance is too great, the joint is dropped. From the remaining joint proposals, the top-k closest to the average are selected, to drop outliers again, and the final joint location is calculated as the average of those. In general, this late fusion step allows for stronger outlier filtering, because multiple proposals can be used, but also works if a person is only visible in two views. In step (12) clearly invalid persons, with too few keypoints, that are too small or large, or outside the room, are dropped, and missing joints are filled with their next neighbors.
@@ 3.3 Performance
Pair filtering (2) 10
Triangulate and Score (3-7) 26
Grouping (9) 4
@@ 4.3 Using more cameras
Because the cameras use fisheye lenses, the images are distorted stronger than in the other datasets. Therefore a different undistortion method is required. Since this requires larger changes in the pipeline of most algorithms, only a subset of the better performing algorithms was tested on this dataset. In some cases the poses can be undistorted after the 2D estimation step, in other cases, the images need to be undistored instead. VoxelPose and Faster-VoxelPose, for which the images are undistorted beforehand, both have problems with a direct transfer to the legoassemble subset, but again show much better performance after synthetic training. The same holds true for VoxelPose on the tennis subset, Faster-VoxelPose on the other hand does not work at all here. The synthetic training always failed with an out-of-memory error, even at the lowest possible batch-size. For mvpose the number of views was too high, the six cameras from tsinghua already took up the complete GPU memory. TEMPO [6] described an evaluation on this dataset, in which they reported an average transfer error of $120\,mm$ (or $36.7\,mm$ after training) on this dataset, but they did not provide further information about the evaluation in the paper or their code. VoxelKeypointFusion shows a very good performance, which is only surpassed by RapidPoseTriangulation.

### v2

@@ 1 Introduction
The ability to accurately detect and track human body positions and movements is an important task in many advanced computer vision applications. From enhancing virtual reality experiences to improving the collaboration with humans in robotic systems, precise human poses play a crucial role at connecting digital systems with the physical world.
@@
However, marker-less pose estimation presents its own set of challenges. Occlusions, varying lighting conditions, and the complexity of human movements can all impact the accuracy and reliability of these systems. To address external and self-occlusion issues, and to some extent also varying lighting, multi-view approaches have become more common. By capturing the scene from multiple angles simultaneously, these systems can overcome many of the limitations that would occur with single-view methods.
Recent advances in deep learning have dramatically improved the accuracy of pose estimation from individual camera views. However, the task of efficiently combining these multiple 2D estimates into accurate 3D poses, particularly for multiple people in real-time scenarios, remains an active area of research. Current approaches often show notable performance drops when applied to setups they were not trained on, and normally are too slow for real-time usage. Furthermore, as applications become more advanced, there is an increasing demand for a more detailed whole-body pose estimation, including facial expressions and especially hand gestures (see Figure 1).
@@ 2 Related Work
Regarding the generalization of learning-based approaches to new datasets, a direct transfer is typically not evaluated. However, VoxelPose, Faster-VoxelPose, MvP, and TEMPO have implemented a finetuning strategy utilizing synthetic data, though only the first two have also published the source-code for this. The idea is to randomly place 3D poses from another dataset in 3D space, and back-project them onto the camera perspectives of the target setup. After adding some augmentations, the network is then trained to learn the 3D reconstruction process. In the closed-source approach of CloseMoCap [16] this concept was further improved by using a larger 3D pose dataset and additional augmentations. Algorithmic approaches generally evaluated their transferability to a limited number of other datasets, a process that is considerably more straightforward as it doesn’t require any training.
@@ 3.1 Algorithm
Similar to most other approaches, the algorithm can be split into two stages as well, with the first one predicting the 2D poses for each image. For this any 2D pose estimator can be integrated, here RTMPose [12] was used. The second stage can be split into the following steps:
@@
In step (1) for each 2D pose, all possible pairs with poses from other views are created, since no association to real persons is known yet. In the next step (2), the pairs are filtered using the previous 3D poses, if available. For this, the 3D poses from the last frame are projected into the 2D views and a pixel-based distance threshold is used. The idea is to reduce the total number of pairs that need to be triangulated. In a pair of 2D persons in two different views, either both parts can match to one of the reprojected 3D poses, only one part is matched, or none of them. If both pairs match, it is likely a valid pair belonging to that 3D person, and it is kept. If only one part matches, a match is unlikely, because the 3D pose should be matchable in both views, so the second part is likely a pose from a different 3D person, and the pair is dropped. If none is matching, it is likely that a new 3D person is present in the current frame, and the pair is kept. In step (3) the core joints are extracted, which are the shoulders, hips, elbows, wrists, knees, and ankles. They are enough to associate the following 3D proposals to a person, and this reduces the computational complexity in the next steps. In step (4) each pair is triangulated to a 3D proposal (see Figure 2(b)). In the following step (5) clearly invalid proposals outside the observed room are dropped. In step (6) the remaining proposals are reprojected into the 2D views. Step (7) calculates a pixel-distance-based reprojection error to the pair’s original 2D poses, and also prunes clearly invalid limbs. In step (8) pairs/proposals with a large error are dropped, as they are considered invalid matches. In step (9) the remaining 3D proposals are grouped in 3D space. Each of the 2D view pairs created a 3D proposal, and if they are close in 3D space, they likely belong to the same person, and are therefore grouped together. Grouping in 3D makes use of physical constraints, which allows a simpler and faster implementation. In step (10) the remaining pairs, which normally are less than in step (4), are triangulated again, but this time with all joints. In step (11) for each group the calculated 3D proposals are merged into a single 3D pose. To filter outliers, first, for each joint an average location is calculated, and then for each proposal, the distance of this joint to the average is calculated. If the distance is too great, the joint is dropped. From the remaining joint proposals, the top-k closest to the average are selected, to drop outliers again, and the final joint location is calculated as the average of those. In general, this late fusion step allows for stronger outlier filtering, because multiple proposals can be used, but also works if a person is only visible in two views. In step (12) clearly invalid persons, with too few keypoints, that are too small or large, or outside the room, are dropped, and missing joints are filled with their next neighbors. The general idea behind the algorithmic structure is to use only the minimal necessary information at each step, drop bad proposals early on, and move expensive operations, like the full-body keypoint calculations, to later steps.
@@ 3.3 Performance
Pair filtering (2) 8
Triangulate and Score (3-7) 27
Grouping (9) 6
@@ 4.3 Using more cameras
Because the cameras use fisheye lenses, the images are distorted stronger than in the other datasets. Therefore a different undistortion method is required. Since this requires larger changes in the pipeline of most algorithms, only a subset of the better performing algorithms was tested on this dataset. In some cases the poses can be undistorted after the 2D estimation step, in other cases, the images need to be undistored instead. VoxelPose and Faster-VoxelPose, for which the images are undistorted beforehand, both have problems with a direct transfer to the legoassemble subset, but again show much better performance after synthetic training. The same holds true for VoxelPose on the tennis subset, Faster-VoxelPose on the other hand does not work at all here. The synthetic training always failed with an out-of-memory error, even at the lowest possible batch-size. For mvpose the number of views was too high, the six cameras from tsinghua already took up the complete GPU memory. TEMPO [6] described an evaluation on this dataset, in which they reported an average transfer error of $120\,mm$ (or $36.7\,mm$ after training) on this dataset, but they did not provide further information about the evaluation in the paper or their code. VoxelKeypointFusion proofs its generalization capabilities with a very good performance, which is only surpassed by RapidPoseTriangulation.